the following series 1-1/2^2-1/3^2+1/4^2-1/5^2-1/6^2+... is convergent or not?
1
Expert's answer
2020-04-15T15:11:30-0400

"1-\frac{1}{2^2}-\frac{1}{3^2}+\frac{1}{4^2}-\frac{1}{5^2}-\frac{1}{6^2}+...(*)"

Consider a series of modules

"1+\frac{1}{2^2}+\frac{1}{3^2}+\frac{1}{4^2}+\frac{1}{5^2}+\frac{1}{6^2}+...=\\
=\sum\limits_{n=1}^{\infty}\frac{1}{n^2}"

A series "\sum\limits_{n=1}^{\infty}\frac{1}{n^{\alpha}}" is convergent if "\alpha>1" ,

 divergent if "\alpha \leq1" .

 In our case "\alpha=2>1" ,

 so the series of modules is convergent, hence the series (*) is

 also convergent.